A group of students were surveyed about the number of siblings they have. Their responses are summarized in the relative frequency table below.

What is the cumulative relative frequency of students with 2 or fewer siblings?

Number of Siblings Relative Frequency
0 0.20
1 0.32
2 0.16
3 0.12
4 or more 0.20
Answer #1

Solution :

From given table,

The cumulative relative frequency of students with 2 or fewer siblings is,

= P(X \leq 2)

= P(X = 0) + (X = 1) + P(X = 2)

= 0.20 + 0.32 + 0.16

= 0.68